# Epistemic Horizons From Deterministic Laws: Lessons From a Nomic Toy Theory

### v1

@@ 1 Introduction
In Newtonian physics one often ignores an explicit account of measurement interactions and that they necessarily disturb the system being measured [2, Chapter 3]. This is traditionally justified by stipulating that the disturbance is determinable and thus can be accounted for. Adjusting one’s measurement record based on known disturbance — if indeed possible — allows an agent to acquire arbitrary information about a system.
Particularly in the context of quantum theory, measurements are said to introduce disturbance. Heisenberg’s uncertainty relations were interpreted by himself as originating from an inevitable measurement disturbance: Whatever pre-determines the outcome of a measurement of a particle is inadvertently disturbed by its interaction with the apparatus [29]. [^2] However, this early account of Heisenberg’s uncertainty is but one possible interpretation. The properties that ‘cause’ an individual measurement result need not exist in a quantum world. In particular, it is unclear whether a single particle can be said to possess properties of position and momentum prior to measurement (see, for example, [20, Section 6.1]). Thus, one cannot straightforwardly argue that such properties (since they do not exist) would be disturbed in a measurement.
@@ 2.2 Properties of Toy Systems
of the set $\mathsf{V}$ of ontic states. Variables that induce the same partition are considered to be equivalent.
Note that Poisson variables are valued in a vector space, whose dimension tells us about the potential number of independent scalar properties it can describe. An important special case is when the dimension is $1$, in which case we speak of a functional $\mathsf{V}\to\mathcal{F}$. Such a linear map automatically satisfies Equation 5.
@@
holds for all $i$ and all $j$. Therefore, Poisson variables precisely correspond to properties which, in Spekkens’ toy theory, are assumed to be knowable about the toy system. While this epistemic horizon is traditionally postulated in Spekkens’ toy theory, we derive it in nomic toy theory.
@@ 2.3 Toy Subjects
And so it may happen that the two perspectives disagree. Even if the omniscient viewpoint contains no fundamental uncertainty about all details of the world, an internal agent could be bound to epistemic limitations. Whether the omniscient view is or is not conceivable, it may be unreachable for any agent as a result of the dynamical constraints of the world in which the agent operates.
To study this tension, let us introduce the notion of agents in nomic toy theory. Note that we do not place any anthropocentric constraints on these, our agents are part of nature in the same way that their objects of study are. Since our agents are decidedly minimal and may not fulfil elaborate requirements for agency [48, 36], we also call them toy subjects.
@@
An example of a toy subject is a simple pointer apparatus. The manifest variable would be the value on a scale or the angle of a pointer needle. Inspired by such example, we call the manifest variable $Q$ of $A$ the position of $A$ and its complementary variable $P$ the momentum of $A$. Even though we do not, in general, assume what type of degrees of freedom the manifest ones are, we label them as ‘positions’ for the sake of simplicity.
@@ 4.2 Summary and Outlook
At first glance, our result may be surprising in light of the claims that Newtonian mechanics should in principle allow for arbitrarily precise measurements of the properties of a classical particle. Bear in mind that Liouville’s theorem in Hamiltonian mechanics implies preservation of phase space volume, but does not rule out arbitrary stretching and squeezing of a phase space volume such that conjugate variables become simultaneously sharply defined. However, we suspect that our result could be related to the claims of de Gosson on the relationship of symplectic geometry and quantum uncertainty principles [17]. Basically, de Gosson derived an analogue of the quantum Robertson–Schrödinger inequality from the symplectic properties of the phase space alone. This essentially implies that Heisenberg’s uncertainty relations already hold in Hamiltonian mechanics for all pairs of conjugate position and momentum variables. Why does it seem that some aspects of quantum uncertainty can be explained in terms of Hamiltonian mechanics? Do uncertainty relations really have such an analogue in classical physics? Can the epistemic horizon in nomic toy theory be restated as a classical uncertainty relation akin to Heisenberg’s uncertainty principle in quantum theory? We hope that our analysis will serve as a toy example to facilitate explorations of those pertinent questions.
@@
A related question concerns the development of ontological models motivated and evaluated from within nomic toy theory. That is, one may investigate what kind of ontologies are consistent with the experience of an epistemically restricted toy subject. Is there a way to differentiate among them based on desiderata such as parsimony or naturalness. In a nutshell, what would such a subject conclude about the ontology underlying the phenomena observed? See also a potential link to problems about bootstrapping and reliabilist epistemology [25].
@@ Appendix A Composing Position and Momentum Measurements
The joint system $A_{1}\oplus S\oplus A_{2}$ starts out at time 00 in the ontic state denoted by
$u(0)=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{3}\\ u_{4}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}$ (28)
@@
The first interaction $m_{p}$ is a measurement of the momentum of $S$ by the toy subject $A_{2}$. Just as at the end of Section 3.1, we substitute the matrix form
@@
$M_{p}=\begin{pmatrix}1&0&0&1\\ 0&1&0&0\\ 0&1&1&0\\ 0&0&0&1\end{pmatrix}$ (30)
Here, we merely write its action on $\mathsf{S}\oplus\mathsf{A}_{2}$. The action on the full joint state space is then via $\mathds{1}_{\mathsf{A}_{2}}\oplus M_{p}$.
At time $1$, i.e. once the interaction $m_{p}$ has taken place, the joint state of all three toy systems is thus
$u(1)=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{3}+u_{6}\\ u_{4}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{4}+u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (31)
As we can see, the manifest variable of $A_{2}$ now encodes the initial momentum of $S$, provided that $A_{2}$ started out in a ready state. Furthermore, the position of $S$ has been disturbed by the initial momentum of $A_{2}$.
The second step of the composite transformation depicted in Figure 3 is a measurement $m_{q}$ of the position of $S$ by the toy subject $A_{1}$. Its matrix form is as in Equation 25 just with the subject and object in reverse order:
$M_{q}=\begin{pmatrix}1&0&1&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&-1&0&1\end{pmatrix}.$ (32)
After this interaction, at time $2$, the full ontic state is
$u(2)=\begin{pmatrix}u_{1}+u_{3}+u_{6}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{3}+u_{6}\\ u_{4}-u_{2}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{4}+u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (33)
@@
$\begin{pmatrix}u_{3}+u_{6}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{3}+u_{6}\\ u_{4}-u_{2}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{4}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (34)
The values of the manifest variables at time $2$ are thus $u_{3}+u_{6}$ and $u_{4}$ respectively. The former encodes the correct position of $S$ at times $1$ and $2$, while the latter encodes the correct momentum of $S$ at times 00 and $1$.
@@ B.1 Systems in Spekkens’ Toy Theory
If the underlying field of $\mathsf{V}$ is that of real numbers, we get continuous toy systems. Basic finite systems are associated with an integer $d$. Their ontic state space is a (symplectic) $\mathbb{Z}_{d}$-module, which is a vector space if $d$ is a prime power. Other finite systems can be obtained as composites of the basic ones (see Section B.2.1).
@@
An epistemic state of Spekkens’ toy theory $(\mathsf{U},a)$ is specified by an isotropic subspace $\mathsf{U}$ of $\mathsf{V}$ and a vector $a\in\mathsf{V}$. Via an isomorphism of $\mathsf{V}$ and its dual $\mathsf{V}^{*}$, the subspace $\mathsf{U}$ is interpreted as consisting of those functionals whose values are known. Alternatively, we can think of $\mathsf{U}$ as the set of values of the orthogonal projection $U\colon\mathsf{V}\to\mathsf{U}$. This is an isotropic variable if and only if $\mathsf{U}$ is isotropic. The vector $a$ is interpreted as one of the ontic states that is deemed possible by this epistemic state. It fixes the value of any functional $u\in\mathsf{U}$ to be
@@ B.2.1 Joint states
Each ontic state of the joint (bipartite) system is given by a pair on ontic states from each of the components respectively. Its underlying vector space is thus the direct product of the individual ones, which is isomorphic to their direct sum.
@@ B.3 General physical transformations
The “if” direction is immediate For the “only if” part, note that since the symplectic form is non-degenerate and the symplectic part of $f$ preserves it, the image of $f$ must coincide with $\mathsf{W}$. Thus, by the first isomorphism theorem, we have
@@ B.5 Pointer-preserving measurements
A specific class of measurements in nomic toy theory are those that preserve pointer states. As we show in Remark B.13 below, they have a special property that contingent manifest variable is trivial. In particular, for a pointer-preserving measurement, the post-measurement manifest variable of $A$ is independent of its the initial momentum, as a result of the property $M_{\mathsf{Q}\mathsf{P}}=0$. For this reason, they are transformations that measure the variable $M_{\mathsf{Q}\mathsf{S}}$ (see Definition 2.9). The characterization of pointer-preserving measurements from Remark B.13 says that they satisfy an additional property, namely that $M_{\mathsf{P}\mathsf{P}}$ is non-degenerate. This suggests that there may be other transformations besides pointer-preserving measurements that also measure the variable $M_{\mathsf{Q}\mathsf{S}}$.
@@
A pointer-preserving merasurement consists of an affine symplectic transformation $m\colon\mathsf{S}\oplus\mathsf{A}\to\mathsf{S}\oplus\mathsf{A}$ and a Lagrangian subspace $\mathsf{Q}$ of $\mathsf{A}$, such that for every $s\in\mathsf{S}$, the associated map $\mathsf{A}\to\mathsf{A}$ given by the composite (here, $A\colon\mathsf{S}\oplus\mathsf{A}\to\mathsf{A}$ denotes the projection map)
@@
Proposition B.12 (Characterization of pointer-preserving measurements).
@@
Let us analyze how a pointer state $(\mathsf{Q},a)$ is transformed by the respective maps in the composite (69). After the first step, adjoining the ontic state $s$ to the possibilistic state $\mathsf{P}+a$ (note that we have $\mathsf{P}=\mathsf{Q}^{T}$ by definition of the momentum variable) leads to the possibilistic state [^10] Note that here we use the same notation for possibilistic states as for epistemic states. That is, for any subspace $\mathsf{U}$ (which need not be isotropic), the possibilistic state associated to $(\mathsf{U},a)$ is $\mathsf{U}^{T}+a$. This notation works for any possibilistic state that is an affine subspace.
@@
Note that if we write $M$ in the block form as in Equation 12, then condition (ii) says that

### v2

@@ 1 Introduction
In Newtonian physics one often ignores an explicit account of measurement interactions and that they necessarily disturb the system being measured [2, Chapter 3]. This is traditionally justified by stipulating that the disturbance is determinable and thus can be accounted for. Adjusting one’s measurement record based on known disturbance — if indeed possible — allows an agent to acquire arbitrary information about a system. Particularly in the context of quantum theory, measurements are said to introduce disturbance. Heisenberg’s uncertainty relations were interpreted by himself as originating from an inevitable measurement disturbance: Whatever pre-determines the outcome of a measurement of a particle is inadvertently disturbed by its interaction with the apparatus [29]. [^2] However, this early account of Heisenberg’s uncertainty is but one possible interpretation. The properties that ‘cause’ an individual measurement result need not exist in a quantum world. In particular, it is unclear whether a single particle can be said to possess properties of position and momentum prior to measurement (see, for example, [20, Section 6.1]). Thus, one cannot straightforwardly argue that such properties (since they do not exist) would be disturbed in a measurement.
@@ 2.1 Toy Systems
A physical state of the toy system (an ontic state) is then specified by an element of $\mathsf{V}$.
@@ 2.2 Properties of Toy Systems
of the set $\mathsf{V}$ of ontic states. Variables that induce the same partition are considered to be equivalent. Note that Poisson variables are valued in a vector space, whose dimension tells us about the potential number of independent scalar properties it can describe. An important special case is when the dimension is $1$, in which case we speak of a functional $\mathsf{V}\to\mathcal{F}$. Such a linear map automatically satisfies Equation 5.
@@
holds for all $i$ and all $j$. Therefore, Poisson variables precisely correspond to properties which, in Spekkens’ toy theory, are assumed to be knowable about the toy system.
While this epistemic horizon is traditionally postulated in Spekkens’ toy theory, we derive it in nomic toy theory.
@@ 2.3 Toy Subjects
And so it may happen that the two perspectives disagree. Even if the omniscient viewpoint contains no fundamental uncertainty about all details of the world, an internal agent could be bound to epistemic limitations. Whether the omniscient view is or is not conceivable, it may be unreachable for any agent as a result of the dynamical constraints of the world in which the agent operates. To study this tension, let us introduce the notion of agents in nomic toy theory. Note that we do not place any anthropocentric constraints on these, our agents are part of nature in the same way that their objects of study are. Since our agents are decidedly minimal and may not fulfil elaborate requirements for agency [48, 36], we also call them toy subjects.
@@
An example of a toy subject is a simple pointer apparatus. The manifest variable would be the value on a scale or the angle of a pointer needle. Inspired by such example, we call the manifest variable $Q$ of $A$ the position of $A$ and its complementary variable $P$ the momentum of $A$.
Even though we do not, in general, assume what type of degrees of freedom the manifest ones are, we label them as ‘positions’ for the sake of simplicity.
@@ 4.2 Summary and Outlook
At first glance, our result may be surprising in light of the claims that Newtonian mechanics should in principle allow for arbitrarily precise measurements of the properties of a classical particle. Bear in mind that Liouville’s theorem in Hamiltonian mechanics implies preservation of phase space volume, but does not rule out arbitrary stretching and squeezing of a phase space volume such that conjugate variables become simultaneously sharply defined. However, we suspect that our result could be related to the claims of de Gosson on the relationship of symplectic geometry and quantum uncertainty principles [17]. Basically, de Gosson derived an analogue of the quantum Robertson–Schrödinger inequality from the symplectic properties of the phase space alone. This essentially implies that Heisenberg’s uncertainty relations already hold in Hamiltonian mechanics for all pairs of conjugate position and momentum variables.
Why does it seem that some aspects of quantum uncertainty can be explained in terms of Hamiltonian mechanics? Do uncertainty relations really have such an analogue in classical physics? Can the epistemic horizon in nomic toy theory be restated as a classical uncertainty relation akin to Heisenberg’s uncertainty principle in quantum theory? We hope that our analysis will serve as a toy example to facilitate explorations of those pertinent questions.
@@
A related question concerns the development of ontological models motivated and evaluated from within nomic toy theory. That is, one may investigate what kind of ontologies are consistent with the experience of an epistemically restricted toy subject. Is there a way to differentiate among them based on desiderata such as parsimony or naturalness? In a nutshell, what would such a subject conclude about the ontology underlying the phenomena observed? See also a potential link to problems about bootstrapping and reliabilist epistemology [25].
@@ Appendix A Composing Position and Momentum Measurements
The joint system $A_{1}\oplus S\oplus A_{2}$ starts out at time $t_{0}$ in the ontic state denoted by
$u(t_{0})=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{3}\\ u_{4}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}$ (28)
@@
The first interaction $m_{p}$ is a measurement of the momentum of $S$ by the toy subject $A_{1}$. Just as at the end of Section 3.1, we substitute the matrix form
@@
$M_{p}=\begin{pmatrix}1&0&0&1\\ 0&1&0&0\\ 0&1&1&0\\ 0&0&0&1\end{pmatrix},$ (30)
where one ought to be careful that the subject and object are now in reverse order compared to Equation 18. Here, we merely write its action on $\mathsf{A}_{1}\oplus\mathsf{S}$. The action on the full joint state space is then via $M_{p}\oplus\mathds{1}_{\mathsf{A}_{2}}$.
At time $t_{1}$, i.e. once the interaction $m_{p}$ has taken place, the joint state of all three toy systems is thus
$u(t_{1})=\begin{pmatrix}u_{1}+u_{4}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{2}+u_{3}\\ u_{4}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (31)
As we can see, the manifest variable of $A_{1}$ now encodes the initial momentum of $S$, provided that $A_{1}$ started out in a ready state. Furthermore, the position of $S$ has been disturbed by the initial momentum of $A_{1}$.
The second step of the composite transformation depicted in Figure 3 is a measurement $m_{q}$ of the position of $S$ by the toy subject $A_{2}$. Its matrix form is as in Equation 25:
$M_{q}=\begin{pmatrix}1&0&0&0\\ 0&1&0&-1\\ 1&0&1&0\\ 0&0&0&1\end{pmatrix}.$ (32)
After this interaction, at time $t_{2}$, the full ontic state is
$u(t_{2})=\begin{pmatrix}u_{1}+u_{4}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{2}+u_{3}\\ u_{4}-u_{6}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{2}+u_{3}+u_{5}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (33)
@@
$\begin{pmatrix}u_{4}\\ u_{2}\end{pmatrix}_{\mathsf{A}_{1}}\oplus\begin{pmatrix}u_{2}+u_{3}\\ u_{4}-u_{6}\end{pmatrix}_{\mathsf{S}}\oplus\begin{pmatrix}u_{2}+u_{3}\\ u_{6}\end{pmatrix}_{\mathsf{A}_{2}}.$ (34)
The values of the manifest variables at time $t_{2}$ are thus $u_{4}$ and $u_{2}+u_{3}$ respectively. The former encodes the correct momentum of $S$ at times $t_{0}$ and $t_{1}$, while the latter encodes the correct position of $S$ at times $t_{1}$ and $t_{2}$.
@@ B.1 Systems in Spekkens’ Toy Theory
If the underlying field of $\mathsf{V}$ is that of real numbers, we obtain continuous toy systems. Basic finite systems are associated with an integer $d$. Their ontic state space is a (symplectic) $\mathbb{Z}_{d}$-module, which is a vector space if $d$ is a prime power. Other finite systems can be obtained as composites of the basic ones (see Section B.2.1).
@@
An epistemic state of Spekkens’ toy theory $(\mathsf{U},a)$ is specified by an isotropic subspace $\mathsf{U}$ of $\mathsf{V}$ and a vector $a\in\mathsf{V}$. Via an isomorphism of $\mathsf{V}$ and its dual $\mathsf{V}^{*}$, the subspace $\mathsf{U}$ is interpreted as consisting of those functionals whose values are known. Alternatively, we can think of $\mathsf{U}$ as the set of values of the orthogonal projection $U\colon\mathsf{V}\to\mathsf{U}$. This is an isotropic variable if and only if $\mathsf{U}$ is isotropic.
The vector $a$ is interpreted as one of the ontic states that is deemed possible by this epistemic state.
It fixes the value of any functional $u\in\mathsf{U}$ to be
@@ B.2.1 Joint states
Each ontic state of the joint (bipartite) system is given by a pair of ontic states from each of the components respectively. Its underlying vector space is thus the direct product of the individual ones, which is isomorphic to their direct sum.
@@ B.3 General physical transformations
The “if” direction is immediate. For the “only if” part, note that since the symplectic form is non-degenerate and the symplectic part of $f$ preserves it, the image of $f$ must coincide with $\mathsf{W}$. Thus, by the first isomorphism theorem, we have
@@ B.5 Pointer-preserving measurements
A specific class of measurements in nomic toy theory are those that preserve pointer states. As we show in Remark B.13 below, they have the special property that the contingent manifest variable is trivial. In particular, for a pointer-preserving measurement, the post-measurement manifest variable of $A$ is independent of its initial momentum, as a result of the property $M_{\mathsf{Q}\mathsf{P}}=0$. For this reason, they are transformations that measure the variable $M_{\mathsf{Q}\mathsf{S}}$ (see Definition 2.9). The characterisation of pointer-preserving measurements from Remark B.13 says that they satisfy an additional property, namely that $M_{\mathsf{P}\mathsf{P}}$ is non-degenerate. This suggests that there may be other transformations besides pointer-preserving measurements that also measure the variable $M_{\mathsf{Q}\mathsf{S}}$.
@@
A pointer-preserving measurement consists of an affine symplectic transformation $m\colon\mathsf{S}\oplus\mathsf{A}\to\mathsf{S}\oplus\mathsf{A}$ and a Lagrangian subspace $\mathsf{Q}$ of $\mathsf{A}$, such that for every $s\in\mathsf{S}$, the associated map $\mathsf{A}\to\mathsf{A}$ given by the composite (here, $A\colon\mathsf{S}\oplus\mathsf{A}\to\mathsf{A}$ denotes the projection map)
@@
Proposition B.12 (Characterisation of pointer-preserving measurements).
@@
Let us analyse how a pointer state $(\mathsf{Q},a)$ is transformed by the respective maps in the composite (69). After the first step, adjoining the ontic state $s$ to the possibilistic state $\mathsf{P}+a$ (note that we have $\mathsf{P}=\mathsf{Q}^{T}$ by definition of the momentum variable) leads to the possibilistic state [^10] Note that here we use the same notation for possibilistic states as for epistemic states. That is, for any subspace $\mathsf{U}$ (which need not be isotropic), the possibilistic state associated to $(\mathsf{U},a)$ is $\mathsf{U}^{T}+a$. This notation works for any possibilistic state that is an affine subspace.
@@
Note that if we write $M$ in block form as in Equation 12, then condition (ii) says that